Calculate the mass of oxalic acid, (COOH)2.2H2O, required to prepare a standard solution with a volume of 100 mL and a concentration of 0.125 mol L-1.
1
Expert's answer
2019-05-01T04:28:02-0400

"m((COOH)2*2H2O)=\dfrac{0.125mol/L *126g/mol *100mL}{1000}=1,575g"